A person standing on the edge of a high cliff throws a rock straight up. The rock misses the edge of the cliff as it fall back to earth in 7s. The rock attains maximum height at t = 1.5 s. What is the height of the cliff? What is the total distance travelled?
1
Expert's answer
2020-09-28T08:06:15-0400

As per the question,

Total time of flight of the rock "(t)=7 sec"

Taken taken by the rock to the reach the maximum height "(t_1)=1.5 sec"

Let the height of the cliff is "h,"

Taken by rock to reach at the bottom from the top of the cliff "(t_2)=7-1.5 = 5.5 sec"

So,

"h= ut + \dfrac{gt_2^2}{2}"

Now, substituting the values in the above equation,

"\Rightarrow h =0+\dfrac{9.8\times 5.5^2}{2}"

"\Rightarrow h=148.225 m"

When the rock will reach to the maximum height, then it's velocity will become zero, let that is represented by "(v)."

Now, "v=u-gt"

Now, substituting the values in the above,

"0=u-9.8\times 1.5 m/sec"

"\Rightarrow u = 14.7 m/sec"

let the distance traveled by stone in 1.5 sec is represented by d,

"v^2=u^2-2gd"

"\Rightarrow d = \dfrac{u^2-v^2}{2g}"

"\Rightarrow d = \dfrac{u^2}{2g}"

"\Rightarrow d = \dfrac{14.7^2}{2\times 9.8}"

"\Rightarrow d = 11.025m"

Hence, total distance traveled by stone "(D)=d+h"

Substituting the values in the above,

"D=(11.025+148.225)m"

"\Rightarrow D=159.25m"